OpenStudy (anonymous):

Find the exact value of sin-1 square root of two divided by two.

OpenStudy (anonymous):

45 degrees

OpenStudy (anonymous):

Just plug \[\sin^{-1} \frac{ \sqrt{2} }{ 2 } \] into your calculator and you should be able to find the exact answer! :)

OpenStudy (anonymous):

It won't let me do that, and 45 isn't one of my answers :(

OpenStudy (anonymous):

is the equation above the right one that you have to solve?

OpenStudy (anonymous):

Oh! it needs it in radians okay so go to this website: http://www.dudefree.com/Student_Tools/materials/precalc/unit-circle.php and you see that next to 45 degrees there is pi/4 so that is your answer :)